Find the angle if its complement is 40 degrees less than one half of its supplement.
A:45 Degrees
B:55 Degrees
C:65 Degrees
D:70 Degrees
E:80 Degrees

Found 2 solutions by macston, Boreal:
Answer by macston(5194)

.
X=the angle; 90-X=complement; 180-X=supplement
.
(90-X)+40=(1/2)(180-X)
130-X=90-(1/2)X Subtract 90 from each side.
40-X=-(1/2)X Add X to each side.
40=(1/2)X Multiply each side by 2.
80=X ANSWER: The angle is 80 degrees.

Answer by Boreal(15235)

x is the angle
complement is (90-x)
supplement is (180-x)
Therefore, (90-x)+40=(1/2)(180-x)
90-x+40=90 -(1/2)x
(x/2)=40
x=80 degrees
complement is 10 degrees
supplement is 100 degrees; half the supplement is 50 degrees, and 40 less is 10 degrees as a check.